# GCRPNet: Graph-Enhanced Contextual and Regional Perception Network for Salient Object Detection in Optical Remote Sensing Images

### v3

@@ Abstract
Salient object detection (SOD) in optical remote sensing images (ORSIs) faces numerous challenges, including significant variations in target scales and low contrast between targets and the background. Existing methods based on vision transformers (ViTs) and convolutional neural networks (CNNs) architectures aim to leverage both global and local features, but the difficulty in effectively integrating these heterogeneous features limits their overall performance. To overcome these limitations, we propose a graph-enhanced contextual and regional perception network (GCRPNet), which builds upon the Mamba architecture to simultaneously capture long-range dependencies and enhance regional feature representation. Specifically, we employ the visual state space (VSS) encoder to extract multi-scale features. To further achieve deep guidance and enhancement of these features, we first design a difference-similarity guided hierarchical graph attention module (DS-HGAM). This module strengthens cross-layer interaction capabilities between features of different scales while enhancing the model’s structural perception, allowing it to distinguish between foreground and background more effectively. Then, we design the LEVSS block as the decoder of GCRPNet. This module integrates our proposed adaptive scanning strategy and multi-granularity collaborative attention enhancement module (MCAEM). It performs adaptive patch scanning on feature maps processed via multi-scale convolutions, thereby capturing rich local region information and enhancing Mamba’s local modeling capability. Extensive experimental results demonstrate that the proposed model achieves state-of-the-art performance, validating its effectiveness and superiority.
Index Terms:
@@
I Introduction
Salient Object detection (SOD) is a fundamental task in the field of computer vision, aiming to identify and segment the most attention-grabbing objects or regions within a scene[1, 2]. As a key prerequisite visual task, research in SOD has been widely applied to areas such as object tracking[3, 4], image segmentation[5, 6], and quality assessment[7, 8]. Traditionally, SOD methods have primarily focused on natural scene images and have achieved remarkable results in that domain. In recent years, with the continuous advancement of satellite remote sensing technology, salient object detection in optical remote sensing images has increasingly attracted the attention of researchers. Compared with natural images, optical remote sensing images pose numerous challenges for salient object detection, such as low target-background contrast, significant variations in target scale, and complex object structures. These factors make it difficult for many SOD methods that perform well on natural images to be directly transferred to remote sensing scenarios. As a result, researchers have developed specialized models tailored to the characteristics of remote sensing images, aiming to effectively detect salient objects in optical remote sensing imagery.
Figure 1: A simple comparison of GCPRNet and other state-of-the-art methods. LiteSalNet[9] and SFANet[10], employing CNNs architectures, produce inaccurate segmentation due to the lack of global context. HFANet[11] adopts a hybrid architecture, but suffers from performance degradation in detection due to the difficulty in integrating the heterogeneous features from CNNs and ViTs.
@@ I Introduction
Deep learning-based ORSI-SOD methods has achieved significant progress over the past decade. The introduction of convolutional neural networks (CNNs) and visual transformers (ViTs), has substantially enhanced the performance of ORSI-SOD models. CNNs, with their local receptive fields, offer high computational efficiency and strong adaptability to scale variations, making them widely adopted in ORSI-SOD tasks. For instance, SFANet[10] enhances multiscale semantic representation through progressive flow and uncertainty-aware refinement. LiteSalNet[9] focuses on efficiency by combining lightweight attention modules with a multistream decoder for accurate saliency detection. Nevertheless, due to the local sliding window operations inherent in CNNs, these methods often lack the ability to model global semantic dependencies. As depicted in LiteSalNet[9] and SFANet[10] from Fig.1, this limitation may lead to segmentation errors and imprecise boundaries, especially when dealing with remote sensing images that contain low-contrast and large-sized objects.
@@
To address this issue, researchers have introduced ViTs into ORSI-SOD tasks to enhance the model’s capacity for global information modeling. ViTs leverage self-attention mechanisms to perform global modeling, effectively compensating for CNNs’ limited receptive field. Their strong global modeling ability makes them particularly advantageous in segmentation tasks involving complex remote sensing imagery. For example, Gao et al.[18] introduced a novel adaptive spatial tokenization Transformer framework that produces high-quality saliency maps. Meanwhile, HFANet[11] integrates a hybrid encoder that combines CNN and Transformer to obtain features containing both global semantic relationships and local spatial details, enabling accurate detection of salient objects in complex backgrounds. However, ViTs lack local inductive bias in modeling spatial structures[19, 20], making it difficult to represent fine-grained object details and often resulting in blurred segmentation boundaries. In hybrid architectures[21, 22, 11], theoretically the local detail modeling of CNNs and the global relational representation of ViTs should be complementary. However, relevant studies[23, 24, 25] have shown that simply combining the two fails to achieve the expected synergistic benefits. This is mainly due to significant differences in their feature representation, receptive field mechanisms, and output paradigm. As depicted in HFANet[11] from Fig.1, these disparities lead to semantic inconsistencies and fusion noise, which in turn hinder effective integration and ultimately degrade overall model performance.
@@
To effectively address the aforementioned challenges, we propose a novel network architecture named graph-enhanced contextual and regional perception network (GCRPNet), which is specifically designed to tackle issues such as low object-background contrast and large variations in object scales, among other issues, in ORSI-SOD tasks. Built upon the conventional encoder-decoder architecture, GCRPNet employs Mamba model[26, 27] as its core modeling backbone. The primary motivation lies in the fact that Mamba employs a selective scanning algorithm to capture long-range dependencies while possessing good stability and scalability, which makes it possible to achieve an effective modeling balance between global context and local region details. To address the challenges of blurred object boundaries and low contrast between salient objects and backgrounds in ORSIs, we propose a difference-similarity guided hierarchical graph attention module (DS-HGAM). This module integrates multi-scale features at skip connections and utilizes a graph neural network[28] to model the spatial relationships and contextual dependencies among the fused features. In this way, our model can better learn the structural differences between foreground and background, thereby reducing segmentation errors caused by boundary ambiguity. Meanwhile, to handle the problem of multi-scale object variations and mitigate the disruption of local region structures caused by Mamba’s scanning process to enhance its local modeling capability, we introduce a locally enhanced visual state space (LEVSS) module. This module first feeds features into a multi-granularity collaborative attention enhancement module (MCAEM) to extract multi-scale local information. It is then followed by a locally enhanced 2D selective scan (LESS2D) operation. Unlike traditional scanning strategies, LESS2D adaptively partitions the feature maps into blocks based on spatial resolution, and performs directional scanning within each block to model the dependencies among local pixels. This method not only strengthens the model’s ability to represent multi-scale features, enhancing its adaptability to scale-diverse targets in ORSIs, but also through a process of block partitioning, directional scanning, and integration, enables the model to enhance local region feature representation while simultaneously achieving global contextual modeling. This improvement in local detail modeling capability further boosts the model’s segmentation accuracy in complex scenes. We conduct extensive comparative experiments on two widely used benchmark datasets (i.e., ORRSD[29] and EORRSD[30]), and the results demonstrate that GCRPNet significantly outperforms existing state-of-the-art methods on the ORSI-SOD task. Our main contributions can be summarized as follows:
@@
    We propose a novel network architecture named GCRPNet, which enhances local region feature representation while achieving global context modeling, and introduces a graph-enhanced module to strengthen multi-scale interactions and boundary modeling, thereby significantly improving the detection accuracy of ORSI-SOD.
@@
    We introduce the DS-HGAM, which leverages a graph attention mechanism to model spatial relationships and contextual dependencies across the fused multiscale features, thus enhancing the model’s structural perception capability and enabling it to effectively learn the differences between foreground and background.
@@
    We propose LEVSS, a module that integrates our proposed MCAEM and LESS2D. It enhances the model’s capability in modeling local region details and effectively combines global context modeling with local feature learning, thereby significantly improving the accuracy and robustness of feature representation.
@@
II RELATED WORK
II-A NSI-SOD
Early SOD methods primarily relied on hand-crafted low-level features and heuristic priors, such as color contrast[15], boundary connectivity[16], and the center-surround assumption. These approaches could achieve reasonable performance in simple backgrounds, but their effectiveness was often limited in complex scenes or multi-object backgrounds due to a lack of semantic understanding of image content. In recent years, deep learning-based SOD methods have gradually become the mainstream[31]. These methods typically adopt end-to-end network architectures that automatically learn saliency representations from large-scale annotated datasets, leading to significant improvements in detection accuracy and fine-grained performance.
CNNs have achieved remarkable progress in the field of SOD due to their powerful multi-level feature learning capabilities. Li et al.[32] significantly improved visual saliency detection performance by leveraging multi-scale features extracted from deep convolutional networks, combined with spatial consistency enhancement and multi-level image segmentation fusion techniques. Sun et al.[33] proposed a novel U-shaped network architecture that effectively integrates multi-scale features and edge information to address issues such as boundary blurring and inaccurate localization commonly found in existing methods. Zhao et al.[34] simultaneously modeled salient object and salient edge information, using edge cues to guide the fusion of salient object features, thereby achieving more accurate localization and boundary detection. With their superior global modeling capabilities and ability to capture long-range dependencies, Transformer-based models have been increasingly applied to NSI-SOD tasks. Liu et al.[35] proposed a pure Transformer-based architecture, rethinking the SOD task from a sequence-to-sequence perspective and enhancing detection performance by modeling global long-range dependencies. Zhang et al.[36] introduced a novel framework based on generative visual Transformers, incorporating latent variables and an energy-based prior distribution to generate uncertainty maps that align with human perception, significantly improving detection performance. Liu et al.[37] proposed TriTransNet, which reconsiders the RGB-D SOD task from the perspective of multi-scale feature fusion. By modeling long-range dependencies across hierarchical features and enhancing RGB representations with depth information, the model achieves more accurate salient object localization and boundary delineation. Recently, Mamba has gained attention in the SOD field due to its strong ability to model long-range dependencies. Zhang et al.[38] introduced a Mamba-based dual-stream feature extractor and a cross-modal fusion module to effectively integrate RGB and depth information, improving salient object detection performance. He et al.[39] proposed a novel unified framework named Samba based on the Mamba architecture, achieving excellent performance and low computational cost across various SOD tasks.
II-B ORSI-SOD
@@
Although SOD has achieved significant progress on natural scene images (NSIs), directly applying existing methods to ORSIs still faces numerous challenges. NSI-SOD methods are primarily designed for natural scenes from a ground-level perspective, typically relying on clear object boundaries, rich textures, and prominent semantic context. However, objects in RSIs often exhibit characteristics such as small size, low contrast, minimal inter-class variance, and strong background interference, making it difficult for existing NSI-based methods to generalize effectively in remote sensing scenarios. To address these issues, a series of SOD approaches specifically tailored for ORSIs—referred to as ORSI-SOD—have emerged in recent years, aiming to enhance detection performance in complex remote sensing environments.
@@
Thanks to their powerful semantic representation and local feature extraction capabilities, CNNs can accurately capture fine-grained local features in ORSIs. In recent years, many CNN-based ORSI-SOD models have achieved notable progress. LV-Net[29] leverages a dual-stream pyramid module and a nested encoder-decoder structure to effectively handle the multi-scale and complex background challenges in ORSIs. DAFNet[30] transmits low-level attention cues to deeper layers to generate accurate high-level attention maps, effectively suppressing background noise and highlighting salient objects. ACCoNet[40] enhances saliency detection by coordinating adjacent features and capturing contextual information, leading to more comprehensive activation of salient regions. MCP-Net[41] introduces a memory module to utilize both inter-image and intra-image contextual information, significantly improving the detection performance for salient objects in optical RSIs. Similarly, Transformer-based methods have been increasingly applied to ORSI-SOD tasks, aiming to overcome the limitations of CNNs in handling complex backgrounds, multi-scale targets, and weak object boundaries in optical remote sensing images. TLCKD-Net[42] adopts a Transformer-based encoder to effectively capture global features, addressing traditional methods’ shortcomings in handling complex backgrounds and salient objects of varying scales and shapes. GeleNet[43] adopts a global-to-local paradigm based on Transformers, combining global feature extraction with local enhancement modules to more precisely locate salient objects and highlight details. However, pure ViT architectures are limited in modeling local information, while targets in optical remote sensing images often exhibit multi-scale and complex shapes, which can easily lead to imprecise boundary segmentation. To address this issue, researchers have proposed hybrid architectures. ASNet[21] integrates Transformer and CNN techniques through a dual-branch encoder that captures both global dependencies and local detail features, adapting to the complexity of RSIs and improving detection performance. ADSTNet[22] combines the strengths of CNNs and Transformers to effectively compensate for global and local information, with boundary-guided modules further assisting in accurate detection and localization.
Figure 2: Pipeline of the proposed method. A: GCRPNet adopts a U-shaped encoder-decoder architecture. The encoder is built upon the visual state space (VSS) of VMamba. In the skip connections, a DS-HGAM module is employed to enhance the model’s representation capability through feature fusion and a graph attention mechanism. Subsequently, in the decoding stage, the features processed by DS-HGAM are fed into the LEVSS blocks for decoding. B: The proposed LEVSS blocks enhance the model’s ability to capture local information through a MCAEM and a LESS2D, aiming for more precise boundary representation. C: LESS2D performs four-directional scanning within each image patch.
II-C Scanning Mechanism in Visual Mamba
Recently, the emergence of state space models (SSMs)[44, 45] has provided a novel paradigm for global modeling, particularly with the introduction of the Mamba[26] model. Mamba enhances global modeling capability by introducing a selective scanning mechanism and adopts a parallel scanning strategy to effectively capture comprehensive contextual information across the image, facilitating better representation of long-range dependencies. Building on this foundation, Mamba has rapidly advanced in the field of computer vision, giving rise to a variety of visual Mamba architectures[46, 47, 48]. Despite Mamba’s impressive potential in vision applications, it still faces challenges in local modeling. In current visual Mamba models[46, 27, 49], a common practice is to divide 2D feature maps into patches and scan them into 1D sequences before feeding them into the SSMs for processing. However, images inherently lack a sequential structure, and this serialization conflicts with the spatial locality of image data. As a result, the dependency relationships between neighboring pixels are disrupted, leading to insufficient modeling of fine-grained structures such as edges and textures. This ultimately hinders the SSMs from generating high-quality visual features.
@@ II-C Scanning Mechanism in Visual Mamba
As a crucial component of Mamba architectures in the vision domain, efficient scanning mechanisms play a key role in improving model performance and have become a focal point of current Mamba-related research. Vision Mamba employs a bidirectional scanning mechanism to simultaneously process forward and backward information. VMamba[46] introduces a cross-scanning mechanism that scans images along four different paths to enable more comprehensive feature extraction, and has seen widespread adoption. The Hilbert scanning strategy[50] integrates multiple scanning methods and directions, allowing it to capture local feature information from various angles. The LocalMamba[51] model incorporates a local scanning strategy by dividing the image into multiple non-overlapping local windows, with pixels in each window scanned independently.
III METHODOLOGY

### v4

# Beyond Global Scanning: Adaptive Visual State Space Modeling for Salient Object Detection in Optical Remote Sensing Images

@@ Abstract
Salient object detection (SOD) in optical remote sensing images (ORSIs) faces numerous challenges, including significant variations in target scales and low contrast between targets and the background. Existing methods based on vision transformers (ViTs) and convolutional neural networks (CNNs) architectures aim to leverage both global and local features, but the difficulty in effectively integrating these heterogeneous features limits their overall performance. To overcome these limitations, we propose an adaptive state space context network (ASCNet), which builds upon the state space model mechanism to simultaneously capture long-range dependencies and enhance regional feature representation. Specifically, we employ the visual state space encoder to extract multi-scale features. To further achieve deep guidance and enhancement of these features, we first design a Multi-Level Context Module (MLCM). This module strengthens cross-layer interaction capabilities between features of different scales while enhancing the model’s structural perception, allowing it to distinguish between foreground and background more effectively. Then, we design the APVSS block as the decoder of ASCNet. This module integrates our proposed Dynamic Adaptive Granularity Scan (DAGS) and Granularity-aware Propagation Module (GPM). It performs adaptive patch scanning on feature maps enhanced by local perception, thereby capturing rich local region information and enhancing state space model’s local modeling capability. Extensive experimental results demonstrate that the proposed model achieves state-of-the-art performance, validating its effectiveness and superiority.
keywords:
@@
journal: Neurocomputing
[…]
[1] organization=School of Computer Science and Technology, Hainan University, city=Haikou, postcode=570228, country=China
[…]
[2] organization=School of Control Science and Engineering, Shandong University, city=Jinan, postcode=250061, country=China
[…]
[3] organization=School of Data Science, Lingnan University, city=Hong Kong, country=China
1 Introduction
Figure 1: Overall performance comparison on the EORSSD(L) and ORSSD(R) datasets. Our method consistently outperforms recent SOTA approaches across multiple evaluation metrics, demonstrating stable and superior performance on both datasets.
Salient Object detection (SOD) is a fundamental task in the field of computer vision, aiming to identify and segment the most attention-grabbing objects or regions within a scene[2, 7]. As a key prerequisite visual task, research in SOD has been widely applied to areas such as object tracking[19, 10], image segmentation[32, 25], and underwater exploration[27], as well as many other fields[48, 33]. Traditionally, SOD methods have primarily focused on natural scene images and have achieved remarkable results in that domain. In recent years, with the continuous advancement of satellite remote sensing technology, salient object detection in optical remote sensing images has increasingly attracted the attention of researchers. Compared with natural images, ORSIs pose numerous challenges for salient object detection, such as low target-background contrast, significant variations in target scale, and complex object structures. These factors make it difficult for many SOD methods that perform well on natural images to be directly transferred to remote sensing scenarios. As a result, researchers have developed specialized models tailored to the characteristics of remote sensing images, aiming to effectively detect salient objects in optical remote sensing imagery.
@@ 1 Introduction
Deep learning-based ORSI-SOD methods have achieved significant progress over the past decade. The introduction of convolutional neural networks (CNNs) and visual transformers (ViTs), has substantially enhanced the performance of ORSI-SOD models. CNNs, with their local receptive fields, offer high computational efficiency and strong adaptability to scale variations, making them widely adopted in ORSI-SOD tasks. For instance, SFANet[34] enhances multiscale semantic representation through progressive flow and uncertainty-aware refinement. RAGRNet[55] enhances saliency detection by integrating CNN features with recurrent graph reasoning to model region–boundary interactions in complex remote sensing scenes. Nevertheless, due to the local sliding window operations inherent in CNNs, these methods often lack the ability to model global semantic dependencies. This limitation may lead to segmentation errors and imprecise boundaries, especially when dealing with remote sensing images that contain low-contrast and large-sized objects.
Figure 2: Qualitative comparison results on challenging ORSIs scenes. From left to right are the RGB images, ground truth (GT), results produced by our method, and those generated by representative state-of-the-art (SOTA) methods. Our approach yields more accurate object localization and boundary delineation, while reducing misdetection and boundary errors.
@@
To address this issue, researchers have introduced ViTs into ORSI-SOD tasks to enhance the model’s capacity for global information modeling. ViTs leverage self-attention mechanisms to perform global modeling, effectively compensating for CNNs’ limited receptive field. Their strong global modeling ability makes them particularly advantageous in segmentation tasks involving complex remote sensing imagery. For example, Gao et al.[11] introduced a novel adaptive spatial tokenization Transformer framework that produces high-quality saliency maps. Meanwhile, HFANet[42] integrates a hybrid encoder that combines CNN and Transformer to obtain features containing both global semantic relationships and local spatial details, enabling accurate detection of salient objects in complex backgrounds. However, ViTs lack local inductive bias in modeling spatial structures[4, 38], making it difficult to represent fine-grained object details and often resulting in blurred segmentation boundaries. In hybrid architectures[46, 54, 42], theoretically the local detail modeling of CNNs and the global relational representation of ViTs should be complementary. However, relevant studies[58, 20, 39] have shown that simply combining the two fails to achieve the expected synergistic benefits. This is mainly due to significant differences in their feature representation, receptive field mechanisms, and output paradigm.
@@
To effectively address the aforementioned challenges, we propose a novel network architecture named adaptive state space context network (ASCNet), which is specifically designed to tackle issues such as low object-background contrast and large variations in object scales in ORSI-SOD tasks. Built upon a conventional encoder–decoder architecture, ASCNet adopts a visual state space model[13, 35] as its core representation framework. The primary motivation lies in the fact that visual state space model employs a selective scanning algorithm to capture long-range dependencies while possessing good stability and scalability, which makes it possible to achieve an effective modeling balance between global context and local region details. To address the challenges of blurred object boundaries and low contrast between salient objects and backgrounds in ORSIs, we propose a Multi-Level Context Module (MLCM). This module integrates multi-scale features across different representation levels and employs a graph neural network[41] to explicitly model spatial relationships and contextual dependencies among regional features. As shown in Figure 2, by capturing cross-regional interactions and structural correlations, MLCM enables the network to better distinguish salient objects from complex backgrounds, thereby reducing segmentation errors caused by boundary ambiguity. Meanwhile, to handle the problem of multi-scale object variations and to alleviate the limitations of sequence modeling in local region representation, we introduce an Adaptive Patchwise Visual State Space (APVSS) block. This module first feeds feature maps into a Granularity-aware Propagation Module (GPM), which enhances local structural perception by introducing global semantic conditioning and promoting consistent interpretation of regional representations. Subsequently, a Dynamic Adaptive Granularity Scan (DAGS) operation is applied to further strengthen local modeling capability. Unlike conventional uniform scan strategies, DAGS partitions feature maps into resolution-aware spatial blocks and performs multi-directional scan within each block, with content-adaptive weighting to modulate the contribution of different directions for capturing local spatial dependencies. Through block-wise partitioning, direction-aware sequence construction, and integration across blocks, DAGS enhances fine-grained local feature representation while preserving global contextual awareness. This direction-adaptive modeling strategy significantly improves the network’s ability to handle scale-diverse targets and complex spatial layouts in ORSIs, ultimately leading to more accurate saliency prediction. We conduct extensive comparative experiments on two widely used benchmark datasets (i.e., ORSSD[23] and EORSSD[52]), and the experimental results demonstrate that ASCNet consistently outperforms existing state-of-the-art methods on ORSI-SOD tasks (as shown in Figure 1). Our main contributions can be summarized as follows:
@@
    We propose a novel network architecture named ASCNet, which enhances local region feature representation while achieving global context modeling, and introduces a topology-aware module to strengthen multi-scale interactions and boundary modeling, thereby significantly improving the detection accuracy of ORSI-SOD.
@@
    We introduce the MLCM, which leverages a topology-aware attention mechanism to model spatial relationships and contextual dependencies across the fused multiscale features, thus enhancing the model’s structural perception capability and enabling it to effectively learn the differences between foreground and background.
@@
    We propose APVSS, a module that integrates our proposed GPM and DAGS. It enhances the model’s capability in modeling local region details and effectively combines global context modeling with local feature learning, thereby significantly improving the accuracy and robustness of feature representation.
@@
2 Related Work
2.1 ORSI-SOD
@@
Although SOD has achieved significant progress on natural scene images (NSIs), directly applying existing methods to ORSIs still faces numerous challenges. NSI-SOD methods [1, 59, 43, 50, 18] are primarily designed for natural scenes from a ground-level perspective, typically relying on clear object boundaries, rich textures, and prominent semantic context. However, objects in RSIs often exhibit characteristics such as small size, low contrast, minimal inter-class variance, and strong background interference, making it difficult for existing NSI-based methods to generalize effectively in remote sensing scenarios. To address these issues, a series of SOD approaches specifically tailored for ORSIs—referred to as ORSI-SOD—have emerged in recent years, aiming to enhance detection performance in complex remote sensing environments.
@@
Thanks to their powerful semantic representation and local feature extraction capabilities, CNNs can accurately capture fine-grained local features in ORSIs. In recent years, many CNN-based ORSI-SOD models have achieved notable progress. ERPNet [57] employs a CNN-based encoder–decoder architecture with edge-guided position attention to progressively refine salient object localization and boundaries in optical remote sensing images. ESGNet [12] exploits edge and skeleton guidance through a two-stage CNN framework, progressively injecting structural cues into saliency features to refine object localization and preserve accurate boundaries in optical remote sensing images. ACCoNet[26] enhances saliency detection by coordinating adjacent features and capturing contextual information, leading to more comprehensive activation of salient regions. Similarly, Transformer-based methods have been increasingly applied to ORSI-SOD tasks, aiming to overcome the limitations of CNNs in handling complex backgrounds and weak object boundaries in optical remote sensing images. TLCKD-Net[8] adopts a Transformer-based encoder to effectively capture global features, addressing traditional methods’ shortcomings in handling complex backgrounds and salient objects of varying scales and shapes. GeleNet[24] adopts a global-to-local paradigm based on Transformers, combining global feature extraction with local enhancement modules to more precisely locate salient objects and highlight details. However, pure ViT architectures are limited in modeling local information, while targets in optical remote sensing images often exhibit multi-scale and complex shapes, which can easily lead to imprecise boundary segmentation. To address this issue, researchers have proposed hybrid architectures. ADSTNet[54] combines the strengths of CNNs and Transformers to effectively compensate for global and local information, with boundary-guided modules further assisting in accurate detection and localization. UDCNet[40] combines CNN-based feature extraction with frequency–spatial Transformer blocks to capture complementary local and global information, while a saliency–edge joint decoder improves precise localization and boundary delineation.
By contrast, our method focuses on adaptive state space contextual modeling, which jointly captures long-range dependencies and fine-grained local structures, effectively addressing boundary ambiguity and scale variation, leading to more accurate saliency localization in complex ORSIs.
2.2 Scan Mechanism in Visual Mamba
@@
Recently, the emergence of state space models (SSMs)[15, 36] has provided a novel paradigm for global modeling, particularly with the introduction of the Mamba[13] model. Mamba enhances global modeling capability by introducing a selective scan mechanism and adopts a parallel scan strategy to effectively capture comprehensive contextual information across the image, facilitating better representation of long-range dependencies. Building on this foundation, Mamba has rapidly advanced in the field of computer vision, giving rise to a variety of visual Mamba architectures[30, 29, 44]. As a crucial component of Mamba architectures in the vision domain, efficient scan mechanisms play a key role in improving model performance and have become a focal point of current Mamba-related research. Vision Mamba employs a bidirectional scan mechanism to simultaneously process forward and backward information. VMamba[30] introduces a cross-scan mechanism that scans images along four different paths to enable more comprehensive feature extraction, and has seen widespread adoption. The Hilbert scan strategy[17] integrates multiple scan methods and directions, allowing it to capture local feature information from various angles. The LocalMamba [21] model incorporates a local scan strategy by dividing the image into multiple fixed-size, non-overlapping local windows, within which pixels are scanned independently.
Figure 3: Pipeline of the proposed method. A: ASCNet adopts encoder-decoder architecture. The encoder is built upon the state space block. In the skip connections, a MLCM is employed to enhance the model’s representation capability through feature fusion and a topology-aware attention mechanism. Subsequently, in the decoding stage, the features processed by MLCM(D) are fed into the APVSS blocks for decoding. B: The proposed APVSS blocks enhance the model’s ability to capture local information through a GPM(E) and a DAGS(C), aiming for more precise boundary representation.
Despite Mamba’s impressive potential in vision applications, it still faces challenges in local modeling. In current visual Mamba models[30, 35, 47], a common practice is to divide 2D feature maps into patches and scan them into 1D sequences before feeding them into the SSMs for processing. However, images inherently lack a sequential structure, and this serialization conflicts with the spatial locality of image data. As a result, the dependency relationships between neighboring pixels are disrupted, leading to insufficient modeling of fine-grained structures such as edges and textures. Moreover, treating all scanning directions equally implicitly assumes isotropic spatial dependencies, which is often violated in ORSIs scenes. These limitations lead to insufficient modeling of fine-grained structures such as edges, textures, and elongated objects, ultimately hindering SSMs from generating high-quality and geometry-consistent visual representations.
3 Methodology